A cargo was dropped down a helicopter which hovered at height of h=125m


When will the cargo reach the land?free fall acceleration is g=10m/s2.Air resistance must be ignored


1
Expert's answer
2022-05-18T08:31:49-0400
"t=\sqrt{2h/g}\\
=\sqrt{2*125/10}=5\rm\: s"